OpenStudy (anonymous):

Five times the difference of one-half of a number and two. How would you write that in an algebraic expression?

OpenStudy (anonymous):

5*(.5-x2)

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

But it's not one of my answers.

OpenStudy (anonymous):

try switching it around

OpenStudy (anonymous):

or look on google lol

OpenStudy (anonymous):

So confused lol

OpenStudy (anonymous):

oh man that stinks lol

OpenStudy (anonymous):

lol tell me about it.

OpenStudy (anonymous):

watch

OpenStudy (anonymous):

Convert everything to decimals: 2.5n = 1.2 Divide both sides by 2.5: n = 1.2 / 2.5 = 0.48 Or work in improper fractions: (5/2)n = 6/5 Multiply both sides by 2/5: n = (6/5)(2/5) = 12/25

OpenStudy (anonymous):

i got that by google lol

OpenStudy (anonymous):

can you reply to my messages please